One printer can do a printing job in 5 hours. Another printer can do the same in 11 hours. How long does it take the two printers to do the job working together?
Just having trouble setting the problem up.

Answer by fractalier(6550)

The set up looks like this
t/5 + t/11 = 1 (1 job)
Now multiply by 55 and get
11t + 5t = 55
16t = 55
t = 55/16 hours